OpenStudy (anonymous):

Given the function g(x) = 4x − 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(−4). The value of g(2) is smaller than the value of g(−4). The value of g(2) is the same as the value of g(−4). The values of g(2) and g(−4) cannot be compared.

OpenStudy (anonymous):

Find g(2) first: g(2) = 4*2 - 5 = 3 and now g(-4) g(-4) = 4*(-4) -5 = -23 So find the relation between g(2) and g(-4) which one is greater??

OpenStudy (anonymous):

g(2)

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yes then find the right choice..

OpenStudy (anonymous):

I am here why are you tagging me??

OpenStudy (anonymous):

If g(2) is greater than what will be the choice??

OpenStudy (anonymous):

so that you knew i was talking to you lol.

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

Right dear..